OpenStudy (anonymous):

I need help with Algebra 1 !!! What are parallel and perpendicular lines? How can you tell the difference between the two, and what are regular slopes? Show examples.

OpenStudy (anonymous):

Parallel Lines - http://www.mathsisfun.com/definitions/parallel-lines.html Perpendicular Lines - http://www.mathsisfun.com/definitions/perpendicular.html I am finding a link for regular slopes. All I am getting is crossword puzzles O_O

OpenStudy (bohotness):

no @OpiGeode this is a homework help study website

OpenStudy (anonymous):

I am looking for someone to help me. I don't know what parallel and perpendicular lines are. Nor how to identify them :/

OpenStudy (daniellelovee):

parallel have the same slope and they look the same and go the same way perpendiculars intersect and form an x form

OpenStudy (daniellelovee):

oh and perpendicular lines have opposite reciprocal slopes means if one slope is 2 the other slope will be -1/2

OpenStudy (anonymous):

So the lines -3 and 3 are parallel?

OpenStudy (daniellelovee):

no because parallel lines have the same everything including slop they are like mirror images

OpenStudy (anonymous):

so -3 and 3 would be neither perp or parallel?

OpenStudy (daniellelovee):

yup unless the 3 was 1/3 which would make it perpendicular

OpenStudy (anonymous):

Perp would be -4 and 1/4 or 10/12 and -1 1/5

OpenStudy (anonymous):

Oh no wait, not the second one

OpenStudy (daniellelovee):

-4 and 1/4 is perpendicular and the other ones dont have the same slope nor are they opposite reciprocals

OpenStudy (anonymous):

and parallel would be like 1/2 and 1/2?

OpenStudy (daniellelovee):

yes because they are the same

OpenStudy (anonymous):

So my homework is asking me to write the equation of a line parallel to y = 1/3x - 2 That would basically be something like: y = 1/3x - 4

OpenStudy (anonymous):

Same slope, just different y intercept?

OpenStudy (daniellelovee):

exactly! :D

OpenStudy (daniellelovee):



OpenStudy (anonymous):

and then it also says to write one perpendicular for the same line: y = 1/3x - 2 That would be..... y = -3x -2 ?

OpenStudy (daniellelovee):

as long as the lines look like mirror images and dont touch they are parallel

OpenStudy (daniellelovee):

YES!

OpenStudy (daniellelovee):

you got it :0

OpenStudy (daniellelovee):

:)

OpenStudy (anonymous):

Umm, I think I am starting to get the concept, but can I just post what I get for some questions and you check them real quick?

OpenStudy (daniellelovee):

sure just tag me